OpenStudy (anonymous):

Participation in the U.S. food-stamp program grew from approximately 17.1 million people in 2000 to approximately 23 million in 2004. a)Calculate the number of participants in 2001. b)Predict the number of participants in 2010. please show work

OpenStudy (gw2011):

23-17.1=5.9 million growth 5.9/4=1.475 million growth per year a) 2001: 17.1+1.475=18.575= approximately to 18.6 million b) 2010: (1.475/year growth)(10 years)=14.75 million increase 17.1 (for the year 2000)+14.75=31.85= approximately 31.9 million

OpenStudy (anonymous):

thank you

OpenStudy (gw2011):

You're welcome